OpenStudy (anonymous):

how do i find the inverse function of f(x) = ln ( x+3) - 4 ?

jimthompson5910 (jim_thompson5910):

f(x) = ln ( x+3) - 4 y = ln ( x+3) - 4 x = ln ( y+3) - 4 Now solve for y

OpenStudy (anonymous):

how can i solve ln(y+3)?

jimthompson5910 (jim_thompson5910):

first add 4 to both sides

OpenStudy (anonymous):

ok,,

jimthompson5910 (jim_thompson5910):

how do you undo the natural log function?

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

how do you undo any log function

OpenStudy (anonymous):

well if 5e''x = 11 is = e''x = 11/5?

jimthompson5910 (jim_thompson5910):

use e^x instead of e"x

OpenStudy (anonymous):

so ln(y+3) is y/lm/3?

jimthompson5910 (jim_thompson5910):

to solve for y in x = ln(y), you raise both sides as exponents with bases of 'e' to get e^x = e^(ln(y)) but that just turns into e^x = y y = e^x

jimthompson5910 (jim_thompson5910):

so here's the complete step by step picture of what's going on \[\Large f(x) = \ln(x+3) - 4\] \[\Large y = \ln(x+3) - 4\] \[\Large x = \ln(y+3) - 4\] \[\Large x+4 = \ln(y+3)\] \[\Large e^{x+4} = y+3\] \[\Large e^{x+4}-3 = y\] \[\Large y = e^{x+4}-3\] \[\Large f^{-1}(x) = e^{x+4}-3\]